# Efficient Reinforcemen Learning via Decoupling Exploration and Utilization

### v3

@@ Abstract
Index Terms:
Reinforcement Learning, Sample Efficiency, Exploration, Ensemble Q-LearningI Introduction
@@
In the realm of Reinforcement Learning for continuous control problems, two roughly antithetical methodologies have traditionally prevailed. On one hand, several researchers endeavor to mitigate overestimation bias, for instance, by adopting the minimum of the estimated values or by employing multiple estimates to formulate an approximate lower bound, thereby manifesting a form of pessimism towards the current value function estimation. Conversely, there exists a school of thought advocating that an inherent optimism in approximation estimations effectively fosters exploration within the environment and action space, thereby augmenting the potential for higher reward acquisition. Both approaches have underpinned the development of state-of-the-art algorithms, as evidenced by the literature [20, 21, 22], suggesting a non-mutually exclusive relationship between these methodologies. Consequently, our study aims to amalgamate these seemingly disparate strategies, thereby harnessing the synergistic benefits of both tactical optimism and tactical pessimism. This integrated perspective leads to the conceptualization of tactical optimism and tactical pessimism in the context of RL. Empirical evidence suggests notable variations in the efficacy of these methods across diverse environments. Our objective is to harmonize these ostensibly conflicting paradigms, postulating that the relative contributions of each component may fluctuate in accordance with the specificities of the task at hand. To this end, we bifurcate the role of the actor in Actor-Critic models, introducing two distinct actors: the ’optimism actor’ for exploration and the ’pessimism actor’ for exploitation. An optimist actor plays a crucial role in facilitating exploration; however, in scenarios characterized by significant estimation errors, the pessimist actor becomes instrumental in stabilizing the learning process. This leads to the formulation of our proposed OPARL(Optimistic and Pessimistic Actor Reinforcement Learning) algorithm, wherein our model accommodates a highly optimistic state overlay during the exploration phase, enhancing the likelihood of discovering high-reward regions. Despite this, such a technique may not always converge to the optimal strategy. Conversely, we employ tactical pessimism during the policy recuperation phase to optimize the exploitation of previously identified high-reward behaviors. The overall concept is summarized in Figure 2.
@@
In our approach, we engage with the environment using a tactically optimistic exploration strategy, while simultaneously employing a tactically pessimistic approach for evaluating all data observed thus far. This tactically pessimistic utilization strategy is designed to maximize long-term rewards, effectively mitigating biases introduced by short-term localized rewards. Therefore, even if the exploration strategy leads to suboptimal behavior in task rewards, the utilization strategy can still recalibrate the exploration strategy to support further exploration. More subtly, this allows the exploration strategy to seek better states and rewards. Ultimately, the exploration strategy can generate superior data, rendering the final performance less sensitive to the balance between intrinsic and extrinsic rewards, which is advantageous for the model to explore more effectively in diverse environments. Our series of experiments have demonstrated that OPARL can balance both pessimistic and optimistic algorithms, offering a wider range of exploration (optimistic exploration methods) and more stable performance (pessimistic utilization methods), achieving the best of both worlds and elevating the technological level for single-agent tasks in challenging environments. Our research validates that conducting exploration and training in an environment inclined towards highly optimistic estimations, particularly characterized by an overestimation of expected rewards, is not only feasible but also effective. Simultaneously, by decoupling and rebalancing exploration and exploitation, superior results are achieved. The experiments indicate that OPARL effectively enhances the model’s generalization ability. Our model demonstrates improved states and rewards in four Mujoco environments and 14 out of 18 DMControl environments, while still maintaining the stability of the original model, The effect is as shown in the Figure 1.
II RELATED WORK
II-A Reinforcement Learning
@@
II-B Sample Efficient Reinforcement Learning
@@
Moreover, Data Augmentation and regularization techniques have proven valuable in enhancing the learning efficiency of existing sample data. These techniques generate additional training data from limited interactions or employ regularization to prevent overfitting. However, the implementation of these techniques often necessitates complex algorithmic design and fine-tuning. Furthermore, their efficacy is typically confined to specific tasks and may not generalize across diverse tasks or environments. In contrast, our proposed algorithm is comparatively straightforward, achieving equivalent or superior performance with fewer interactions. It rapidly learns from limited experiments, circumventing the high costs associated with data collection and potential safety risks.
II-C Optimistic and Pessimistic in Reinforcement Learning
@@
However, if highly random regions also possess lower values, overestimation bias might lead to agents excessively exploring low-value areas. Thus, another group of scholars emphasizes the necessity to standardize highly uncertain state-action pairs. Based on the DDPG[40] algorithm, the TD3[19] method selects the smaller of two values[41] as an approximate lower bound and explores the environment pessimistically. By reducing pessimistic exploration in value estimation, agents can achieve better long-term rewards. Therefore, our algorithm maintains this pessimistic approach in utilization strategies to minimize extrapolation errors, ensuring model stability and performance.
III PROBLEM SETUP
@@
To learn behavior that maximizes rewards, we first need to obtain an experience pool buffer containing high-reward regions. The buffer is }, which is collected using the tactically optimistic behavior strategy . The optimistic action is obtained through , and a new state is obtained by interacting with the environment. The reward and whether the action is completed are added to the experience pool to obtain more high-reward regions. When using data from the experience pool, we extract states , actions , new states , rewards , and whether the action is completed from the experience pool and use tactical pessimism to obtain more reasonable actions, thereby maximizing the reward. The specific algorithms of both will be presented in detail in the fourth part.
IV Optimistic and Pessimistic Actors in RL
The aim of this section is to establish a novel algorithm to decouple two distinct functions of actors. The first function involves optimistic experience collection and strategizing to explore areas with higher returns. The second function is designed to penalize outlier state-action pairs, employing tactical pessimism to secure more reasonable actions in high-return areas, thus maximizing rewards. This decoupling allows us to explore the environment optimistically while mitigating biases in overestimating expected rewards in our evaluation strategy. Most online reinforcement learning methods implicitly integrate some level of noise into their exploration strategies to expand their exploration scope, as exemplified by the -greedy noise in DQN or noise in TD3. However, these methods often fall short of enabling comprehensive exploration. Our framework, OPARL, amalgamates the values of multiple state-action pairs to access the broadest possible exploration area.
In Section A, we introduce the general OPARL framework, which can be effectively combined with any reinforcement learning algorithm for enhancement. In Section B, we delve into a mathematical exploration of merging optimistic exploration with pessimistic exploitation, illustrating how this strategy enables an agent to attain high rewards in the face of challenging exploration tasks. Section C explores the implementation of combining optimistic exploration and pessimistic exploitation through parameter resetting, demonstrating the synergy of these approaches.
IV-A The OPARL Framework
Our algorithm is bifurcated into two distinct phases: optimistic exploration and pessimistic utilization. In the exploration phase, optimistic strategic exploration is conducted at every step, determining the action with the maximal Q-value through a post-optimistic strategy state. The action associated with the highest Q-value is selected, signifying the culmination of the optimistic exploration process. Rewards and next states are procured from environmental interactions and are systematically stored in a buffer. In contrast, during the pessimistic phase, actions correlated with the minimal Q-value, as dictated by the pessimistic strategy, are chosen to accrue rewards and next states, which are then amalgamated into the buffer. This facilitates agents in exploring regions with augmented rewards, attributable to an elevated exploration ratio. In the utilization phase, a pessimistic approach is adopted. The state, as ascertained through the pessimistic strategy function, identifies the action with the lowest Q-value for the purposes of loss computation and parameter updating. Parameters from the optimistic strategy are periodically transferred to the pessimistic strategy function at every step, fostering a more profound integration of both optimistic and pessimistic methodologies.The pseudocode for OPARL is shown in Algorithm 1.
IV-B The Algorithm of OPARL
@@
And Formulas 3 stands for random Walkely determined values from 1 to , respectively. From this, we can discern that Formula 3, which involves random selection of state-action pairs among N -values, encapsulates exploration bounds that correspond to optimistic and pessimistic exploration at its upper and lower limits, respectively. This delineates the exploratory scope within which the random strategy operates, framed by the optimistic exploration’s propensity for high-reward regions and the cautious approach characteristic of pessimistic exploration.
@@ IV-C Parameter Reset
V EXPERIMENTS
Our experiments demonstrate the efficacy of decoupling exploration and exploitation strategies across a spectrum of tasks with varying demands for exploration. In Section A, we elucidate the baseline and environmental prerequisites. Section B articulates a comparative analysis of fundamental online reinforcement learning algorithms such as TD3, Soft Actor-Critic(SAC), and Proximal Policy Optimization(PPO)[45] within the Mujoco environment[46][47], juxtaposed against our proposed model. Section C escalates the comparison to contemporary advanced exploration models like Random Network Distillation(RND)[48] and exploration-by-disagreement(EBD)[49], underscoring the superior exploratory prowess of our model, OPARL. In Section D, the complexity is furthered by evaluating OPARL in the more challenging DMControl environment[50], benchmarking against baseline models including TD3 and SAC, as well as the avant-garde model REDQ, accompanied by a thorough analysis. Section E delves into a series of ablation studies designed to substantiate the merits of decoupled exploration and exploitation. These studies dissect the performance of OPARL under scenarios of unmitigated optimism, unrelenting pessimism, and in the absence of an ensemble Q function.
@@ V EXPERIMENTS
BallInCup |
Catch |
983.8 3.8 |
978.8 3.7 |
980.3 3.4 |
979.2 1.0 |
Cartpole |
Balance |
999.8 0.0 |
984.0 6.0 |
997.7 1.5 |
999.8 0.0 |
Cartpole |
BalanceSparse |
1000.0 0.0 |
872.1 262.7 |
997.6 5.7 |
1000.0 0.0 |
Cartpole |
Swingup |
874.6 7.5 |
828.1 17.2 |
865.1 1.6 |
865.9 0.8 |
Cheetah |
Run |
861.6 21.4 |
614.2 58.2 |
873.4 21.5 |
788.7 50.9 |
Finger |
Spin |
974.6 15.5 |
940.1 33.5 |
966.3 27.1 |
949.0 16.2 |
Fish |
Swim |
356.0 65.1 |
159.3 100.1 |
342.4 134.5 |
367.4 59.8 |
Fish |
Upright |
926.0 18.6 |
799.6 113.8 |
898.6 50.4 |
913.0 14.6 |
Hopper |
Stand |
599.5 202.2 |
393.5 225.8 |
597.8 308.8 |
354.4 149.9 |
Pendulum |
Swingup |
424.4 155.0 |
382.6 297.0 |
226.2 228.9 |
382.6 209.3 |
PointMass |
Easy |
909.4 8.5 |
880.9 16.7 |
889.9 33.1 |
892.5 5.2 |
Reacher |
Easy |
983.6 1.2 |
970.9 24.4 |
983.5 4.2 |
983.8 1.2 |
Reacher |
Hard |
977.5 2.9 |
964.1 24.0 |
958.6 40.9 |
934.1 71.2 |
Swimmer |
Swimmer6 |
471.4 60.8 |
215.8 119.0 |
359.3 130.9 |
275.6 84.9 |
Swimmer |
Swimmer15 |
442.1 187.5 |
178.6 116.6 |
264.6 136.9 |
220.4 82.3 |
Walker |
Run |
752.1 23.2 |
590.9 51.6 |
773.0 32.9 |
649.3 82.9 |
Walker |
Stand |
989.0 2.0 |
974.0 12.6 |
986.8 2.7 |
984.5 1.3 |
Walker |
Walk |
975.3 2.9 |
957.3 10.6 |
973.5 4.4 |
969.2 2.1 |
| Average | Scores | 805.7 |
704.7 |
774.1 |
750.5 |
| Improvements | Percentage |
14.30% |
4.08% |
7.36% |
@@ V-A Evaluation Setting
In our research, we subjected the newly proposed OPARL algorithm to benchmark testing, assessing its efficacy relative to a cohort of acknowledged algorithms. Owing to their demonstrated reliability in various tasks, TD3[19], SAC[51], and Proximal Policy Optimization (PPO) [45]were chosen as our principal benchmarks. In exploration comparisons, we selected state-of-the-art models such as RND[48] and EBD[49] for juxtaposition. Additionally, within the DMControl suite, we integrated comparisons with Randomized Ensembled Double Q-Learning (REDQ)[52], which is distinguished by its ensemble Q-technology that augments exploratory capabilities and stability. To guarantee equitable and uniform comparisons, we utilized implementations from CleanRL with default hyperparameters as stipulated by the original authors, and the algorithms of the comparative models are delineated in the appendix C.
@@
Our study utilized the state-based DMControl suite[50] under the Mujoco[46][47] framework provided by OpenAI’s Gym[53], to assess performance across four continuous control tasks. The choice of OpenAI Gym[53] was due to its array of environments conducive to benchmarking reinforcement learning algorithms, enabling efficient algorithm-environment interactions. In the Mujoco environment, each algorithm was evaluated over a million time steps, with five seeds randomly selected every 5000 steps to calculate the average rewards and their variance, indicating the range of performance variability. Consistent with related studies, in the DMControl environment, we randomly selected ten seeds every 10000 steps for calculating average rewards, using the variance from these ten seeds to represent the fluctuation range of the model’s performance. For more environment details, please refer to the appendix B.
@@
We modulate the levels of optimism and pessimism in our model via the ensemble , initially set at five. The state-action pairs are determined by selecting the one with the greatest variance from these five values. To balance optimistic and pessimistic explorations, we set their ratio at 1:1. Additionally, we revise the pessimistic exploration parameters every 20000 time steps and incorporate them into the optimistic exploration parameters . The other parameter settings of this algorithm are in line with the original TD3 model[19]. We believe that adhering to these baseline parameters facilitates a fair comparison with existing methodologies while allowing the nuances of our approach to be highlighted. Further details about the implementation are available in the appendix D.
V-B Evaluation of OPARL in Mujoco Environment
@@
In the Mujoco environment, our results, as depicted in the accompanying figure 4, reveal that the OPARL model exhibits exceptional performance, particularly in the Hopper and Ant scenarios. Notably, OPARL consistently surpasses its foundational algorithm, TD3, in the majority of test cases. A trend observed is the progressive stabilization of the model with increasing time steps, marked by OPARL exhibiting the least fluctuation amplitude among all models compared. This attribute underscores OPARL’s capacity to enhance learning stability. In a comprehensive assessment, particularly in the Ant environment, our model significantly outperforms competing models. Quantitatively, OPARL achieves performance improvements of 44.04%, 51.25% and 773.02% over the SAC, TD3, and PPO[45] models, respectively. At the same time, our algorithm achieves performance enhancements of 50.93% and 497.62% over the RND[48] and EBD[49]. This significant improvement underscores the efficacy of integrating multiple Q-values and adopting a maximal value approach (optimistic exploration strategy) in the decision-making process, which markedly augments the algorithm’s capacity for exploration. Consequently, this approach substantially broadens its adaptability and learning scope within complex environments. The findings affirm the premise that the integration of diverse predictive models can facilitate more extensive and effective environmental exploration, which is crucial for learning in high-dimensional and intricate tasks.
V-C Evaluation of OPARL in DMControl Environment

### v4

# Efficient Reinforcement Learning via Decoupling Exploration and Utilization

@@ Abstract
Keywords:
Reinforcement Learning Sample Efficiency Exploration Ensemble Q-Learning.1 Introduction
@@
In the realm of Reinforcement Learning for continuous control problems, two roughly antithetical methodologies have traditionally prevailed. On one hand, several researchers endeavor to mitigate overestimation bias, for instance, by adopting the minimum of the estimated values or by employing multiple estimates to formulate an approximate lower bound, thereby manifesting a form of pessimism towards the current value function estimation. Conversely, there exists a school of thought advocating that an inherent optimism in approximation estimations effectively fosters exploration within the environment and action space, thereby augmenting the potential for higher reward acquisition. Both approaches have underpinned the development of state-of-the-art algorithms, as evidenced by the literature [20, 21, 22], suggesting a non-mutually exclusive relationship between these methodologies. Consequently, our study aims to amalgamate these seemingly disparate strategies. To this end, we introduce the conceptualization of the OPARL (Optimistic and Pessimistic Actor Reinforcement Learning) algorithm, which bifurcates the role of the actor in the Actor-Critic model into two distinct components: one dedicated to optimistic exploration and the other to pessimistic exploitation. During the exploration phase, our model employs a highly optimistic state overlay, thereby enhancing the probability of identifying regions of high reward. Nonetheless, such a technique may not consistently converge to the optimal strategy. In contrast, during the policy recuperation phase, we implement strategic pessimism to optimize the exploitation of previously recognized high-reward behaviors. The overall concept is summarized in Figure 1(b).
@@
Our series of experiments have demonstrated that OPARL can balance both pessimistic and optimistic algorithms, offering a wider range of exploration and more stable performance. Our research validates that conducting exploration and training in an environment inclined towards highly optimistic estimations, particularly characterized by an overestimation of expected rewards, is not only feasible but also effective. Simultaneously, by decoupling and rebalancing exploration and exploitation, superior results are achieved. The experiments indicate that OPARL effectively enhances the model’s generalization ability. Our model demonstrates improved states and rewards in four Mujoco environments and 14 out of 18 DMControl environment, The effect is as shown in the Figure 1(a).
2 RELATED WORK
2.1 Reinforcement Learning
@@
2.2 Sample Efficient Reinforcement Learning
@@
However, the implementation of these techniques often necessitates complex algorithmic design and fine-tuning. Furthermore, their efficacy is typically confined to specific tasks and may not generalize across diverse tasks or environments. In contrast, our proposed algorithm is comparatively straightforward, achieving equivalent or superior performance with fewer interactions. It rapidly learns from limited experiments, circumventing the high costs associated with data collection and potential safety risks.
2.3 Optimistic and Pessimistic in Reinforcement Learning
@@
However, if highly random regions also possess lower values, overestimation bias might lead to agents excessively exploring low-value areas. Thus, another group of scholars emphasizes the necessity to standardize highly uncertain state-action pairs. Based on the DDPG[40] algorithm, the TD3[19] method selects the smaller of two values[41][42] as an approximate lower bound and explores the environment pessimistically. By reducing pessimistic exploration in value estimation, agents can achieve better long-term rewards. Therefore, our algorithm maintains this pessimistic approach in utilization strategies to minimize extrapolation errors, ensuring model stability and performance.
3 PROBLEM SETUP
@@
To learn behavior that maximizes rewards, we first need to obtain an experience pool buffer containing high-reward regions. The buffer is }, which is collected using the tactically optimistic behavior strategy . The optimistic action is obtained through , and a new state is obtained by interacting with the environment. The reward and whether the action is completed are added to the experience pool to obtain more high-reward regions. When using data from the experience pool, we extract states , actions , new states , rewards , and whether the action is completed from the experience pool and use tactical pessimism to obtain more reasonable actions, thereby maximizing the reward.
4 Optimistic and Pessimistic Actors in RL
In section 4.1, we introduce the general OPARL framework, which can be effectively combined with any reinforcement learning algorithm for enhancement. In section 4.2, we delve into a mathematical exploration of merging optimistic exploration with pessimistic exploitation, illustrating how this strategy enables an agent to attain high rewards in the face of challenging exploration tasks. In section 4.3 explores the implementation of combining optimistic exploration and pessimistic exploitation through parameter resetting, demonstrating the synergy of these approaches.
4.1 The OPARL Framework
Our algorithm is divided into two distinct phases: optimistic exploration and pessimistic utilization. During the exploration phase, state-action pairs with the maximum Q-value are determined through an optimistic strategy, and rewards and subsequent states obtained from interactions with the environment are stored in a buffer. Conversely, during the evaluation phase, we select state-action pairs with the minimum Q-value to compute losses and update parameters. Parameters from the optimistic strategy are periodically transferred to the pessimistic strategy function every steps, fostering a deeper integration of optimistic and pessimistic approaches. The pseudocode for OPARL is illustrated in Algorithm 1.
4.2 The Algorithm of OPARL
@@
And Formula 3 stands for random walk determined values from 1 to , respectively. From this, we can discern that Formula 3, which involves random selection of state-action pairs among N -values, encapsulates exploration bounds that correspond to optimistic and pessimistic exploration at its upper and lower limits, respectively. This delineates the exploratory scope within which the random strategy operates, framed by the optimistic exploration’s propensity for high-reward regions and the cautious approach characteristic of pessimistic exploration.
@@ 4.3 Parameter Reset
5 EXPERIMENTS
Our experiments demonstrate the efficacy of decoupling exploration and exploitation strategies across a spectrum of tasks with varying demands for exploration. In Section 5.1, we elucidate the baseline and environmental prerequisites. In Section 5.2 provides a detailed comparative analysis of a range of online reinforcement learning algorithms within the Mujoco environment, juxtaposing their performance against our proposed model. In Section 5.3, the complexity is heightened by assessing our proposed OPARL algorithm within the challenging DMControl environment[46], benchmarking it against baseline models, and providing an exhaustive analysis. In Section 5.4 delves into a series of ablation studies designed to substantiate the merits of decoupled exploration and exploitation.
@@ 5 EXPERIMENTS
BallInCup | Catch | 983.8 3.8 | 978.8 3.7 | 980.3 3.4 | 979.2 1.0 |
Cartpole | Balance | 999.8 0.0 | 984.0 6.0 | 997.7 1.5 | 999.8 0.0 |
Cartpole | BalanceSparse | 1000.0 0.0 | 872.1 262.7 | 997.6 5.7 | 1000.0 0.0 |
Cartpole | Swingup | 874.6 7.5 | 828.1 17.2 | 865.1 1.6 | 865.9 0.8 |
Cheetah | Run | 861.6 21.4 | 614.2 58.2 | 873.4 21.5 | 788.7 50.9 |
Finger | Spin | 974.6 15.5 | 940.1 33.5 | 966.3 27.1 | 949.0 16.2 |
Fish | Swim | 356.0 65.1 | 159.3 100.1 | 342.4 134.5 | 367.4 59.8 |
Fish | Upright | 926.0 18.6 | 799.6 113.8 | 898.6 50.4 | 913.0 14.6 |
Hopper | Stand | 599.5 202.2 | 393.5 225.8 | 597.8 308.8 | 354.4 149.9 |
Pendulum | Swingup | 424.4 155.0 | 382.6 297.0 | 226.2 228.9 | 382.6 209.3 |
PointMass | Easy | 909.4 8.5 | 880.9 16.7 | 889.9 33.1 | 892.5 5.2 |
Reacher | Easy | 983.6 1.2 | 970.9 24.4 | 983.5 4.2 | 983.8 1.2 |
Reacher | Hard | 977.5 2.9 | 964.1 24.0 | 958.6 40.9 | 934.1 71.2 |
Swimmer | Swimmer6 | 471.4 60.8 | 215.8 119.0 | 359.3 130.9 | 275.6 84.9 |
Swimmer | Swimmer15 | 442.1 187.5 | 178.6 116.6 | 264.6 136.9 | 220.4 82.3 |
Walker | Run | 752.1 23.2 | 590.9 51.6 | 773.0 32.9 | 649.3 82.9 |
Walker | Stand | 989.0 2.0 | 974.0 12.6 | 986.8 2.7 | 984.5 1.3 |
Walker | Walk | 975.3 2.9 | 957.3 10.6 | 973.5 4.4 | 969.2 2.1 |
Average | Scores | 805.7 | 704.7 | 774.1 | 750.5 |
Improvements | Percentage | / | 14.30% | 4.08% | 7.36% |
@@ 5.1 Evaluation Setting
In our research, we subjected the newly proposed OPARL algorithm to benchmark testing, assessing its efficacy relative to a cohort of acknowledged algorithms. Owing to their demonstrated reliability in various tasks, TD3[19], SAC[47], and Proximal Policy Optimization (PPO) [48]were chosen as our principal benchmarks. In exploration comparisons, we selected state-of-the-art models such as RND[49] and EBD[50] for juxtaposition. Additionally, within the DMControl suite, we integrated comparisons with Randomized Ensembled Double Q-Learning (REDQ)[51], which is distinguished by its ensemble Q-technology that augments exploratory capabilities and stability. To guarantee equitable and uniform comparisons, we utilized implementations from CleanRL with default hyperparameters as stipulated by the original authors.
@@
Our study utilized the state-based DMControl suite[46] under the Mujoco[52][53] framework provided by OpenAI’s Gym[54], to assess performance across four continuous control tasks. The choice of OpenAI Gym[54] was due to its array of environments conducive to benchmarking reinforcement learning algorithms, enabling efficient algorithm-environment interactions.
@@
We modulate the levels of optimism and pessimism in our model via the ensemble , initially set at five. The state-action pairs are determined by selecting the one with the greatest variance from these five -values. To balance optimistic and pessimistic explorations, we set their ratio at 1:1. Additionally, we revise the pessimistic exploration parameters every 20000 time steps and incorporate them into the optimistic exploration parameters . The other parameter settings of this algorithm are in line with the original TD3 model[19]. We believe that adhering to these baseline parameters facilitates a fair comparison with existing methodologies while allowing the nuances of our approach to be highlighted.
5.2 Evaluation of OPARL in Mujoco Environment
@@
In the Mujoco environment, our results as depicted in the accompanying Figure 2, reveal that the OPARL model exhibits exceptional performance, particularly in the Hopper and Ant scenarios. Notably, OPARL consistently surpasses its foundational algorithm, TD3, in the majority of test cases. A trend observed is the progressive stabilization of the model with increasing time steps, marked by OPARL exhibiting the least fluctuation amplitude among all models compared. This attribute underscores OPARL’s capacity to enhance learning stability. In a comprehensive assessment, particularly in the Ant environment, our model significantly outperforms competing models. Quantitatively, OPARL achieves performance improvements of 44.04%, 51.25% and 773.02% over the SAC, TD3, and PPO[48] models, respectively. At the same time, our algorithm achieves performance enhancements of 50.93% and 497.62% over the RND[49] and EBD[50]. This significant improvement underscores the efficacy of integrating multiple Q-values and adopting a maximal value approach (optimistic exploration strategy) in the decision-making process, which markedly augments the algorithm’s capacity for exploration. Consequently, this approach substantially broadens its adaptability and learning scope within complex environments.
5.3 Evaluation of OPARL in DMControl Environment